Form the partial differential equation by eliminating the arbitrary function f from z=f(x2-y2)

Form the partial differential equation by eliminating the arbitrary function f from z=f(xy)

Form the partial differential equation by eliminating the arbitrary function from z2-xy=f(x/z)


1
Expert's answer
2021-08-24T02:54:46-0400

1)

"Z=f(x^2-y^2)\\
\dfrac{\delta{Z}}{\delta{x}}=2xf'(x^2-y^2)\\
\dfrac{\delta{Z}}{\delta{y}}=-2yf'(x^2-y^2)\\
\therefore\\
y\dfrac{\delta{Z}}
{\delta{x}}+x\dfrac{\delta{Z}}{\delta{y}}=0"

2)

"z=f(xy)\\
\dfrac{\delta{Z}}{\delta{x}}=yf'(xy)\\
\dfrac{\delta{Z}}{\delta{y}}=xf'(xy)\\
\therefore\\
x\dfrac{\delta{Z}}{\delta{x}}-y\dfrac{\delta{Z}}{\delta{y}}=0"

3)

"z^2-xy=f(\frac{x}{z})\\
2z\dfrac{\delta{z}}{\delta{x}}-y=(\dfrac{1}{z}-\dfrac{x}{z^2}\dfrac{\delta{z}}{\delta{x}} )f'(\frac{x}{z})\\
2z\dfrac{\delta{z}}{\delta{y}}-x=(-\dfrac{x}{z^2}\dfrac{\delta{z}}{\delta{y}} )f'(\frac{x}{z})\\
\therefore\\
\dfrac{\delta{z}}{\delta{y}}(\dfrac{xy}{z^2}-1)-\dfrac{x^2}{z^2}\dfrac{\delta{z}}{\delta{x}}=-\dfrac{x}{z}"